Find the joint probability of A and B if
P(B)=0.60 and P(AIB)=0.70
[Here P(AIB) means probability of A given B,give your answer to nearest 2 dp.]

Answer by Theo(13342)

p(a given b) = p(a and b) / p(b)
p(a given b) = .70
p(b) = .60
find p(a and b)
p(a given b) = p(a and b) / p(b) becomes:
.7 = p(a and b) / .6
solve for p(a and b) to get:
p(a and b) = .7 * .6 = .42
that's your solution.